alexismaldo3030:

A segment with endpoints A (4, 2) and C (1, 5) is partitioned by a point B such that AB and BC form a 1:3 ratio. Find B. (1, 2.5) (2.5, 3.5) (3.25, 2.75) (3.75, 4.5)

Narad:

The points are A=(4,2) C=(1,5) Let point B=(x,y) Then, the x-coordinates are in the ratio 1:3 AB/BC =(x-4)/(1-x)=1/3 Solving this equation in x 3x-12=1-x 4x=13 x=13/4=3.25 The y-coordinates are in the ratio 1:3 AB/BC =(y-2)/(5-y)=1/3 Solving this equation in y 3y-6=5-y 4y=11 y=11/4=2.75 The point B is (3.25, 2.75) This is option C